Assume that in a marketing study it has been found that on a fall Saturday afternoon 20% of adult men watch sports programming on network A from 1:00pm to 2:00pm, 25% watch on network C from 2:00pm to 3:00pm, and 10% watch on both networks. For simplicity, assume that viewers of these programs are sure to see advertisements if they are watching a program. If Chicken Cluck's Restaurants advertise only on network C during the sports programming segment, (a) what is the probability that the ad will be seen by a randomly chosen adult man? (b) what would be the probability if they advertise on both networks? (c) If they advertise on both networks, what is the probability that an ad will be seen only on network A?
Answer #1

Answer:

P(A) = 0.20

P(B) = 0.25

P(A and B) = 0.10

a)

Given,

To determine the probability that the ad will be seen by a randomly chosen adult man

i.e.,

probability that he watches C

= 0.25

b)

To determine the probability if they advertise on both networks

Probability P(A or B) = P(A) + P(B) - P(A and B)

substitute values

= 0.2 + 0.25 - 0.1

P(A or B) = 0.35

Hence the probability if they advertise on both networks is 0.35

c)

To determine the probability that an ad will be seen only on network A

P(Ad will be seen only on A | Advertise on both networks) = P(A) / P(A or B)

= 0.20 / 0.35

= 0.5714

Required probability = 0.5714